# Full-Duplex Millimeter Wave MIMO Channel Estimation: A Neural Network Approach

### v1

@@ I INTRODUCTION
To realize full-duplex transmission, the self-interference (SI) caused by each transmit element over the entire receive array has to be canceled out. Consecutive cancellation procedures are usually considered in three domains: propagation, analog, and digital [4]. In propagation domain cancellation, SI power is suppressed before reaching the receive chain circuit. Analog domain cancellation is performed before the analog-to-digital converter (ADC) in the receiver analog chain. Finally, digital domain cancellation suppresses the SI power remaining from the propagation and analog domains [5]. There are two main antenna configurations used for full-duplex systems: separate and shared. In the former, separate antennas are employed for transmission and reception, yielding high isolation and relatively low interference. In contrast, in the latter configuration, a single antenna is utilized for both functions, offering a simpler design and lower hardware costs [6].
@@
As highlighted in several papers, such as [2], [3], and [13], the channel estimation problems in full-duplex transmission have not received extensive research attention in the literature. The few existing works mentioned above are primarily focused on simple channel estimation techniques, like LS estimators, which may not ensure high-quality channel estimates in low signal-to-noise ratios (SNRs). On the other hand, the MMSE channel estimation requires knowledge of the channel correlation matrix and is burdened with high computational complexity. Therefore, it is crucial to thoroughly investigate the development of low-complexity and precise channel estimators with low pilot overhead for full-duplex systems. Furthermore, in separate antenna configurations, where the receive antenna (RX) chains lack the capability to receive pilot signals, the channel estimation from transmit antenna (TX) arrays to the downlink UEs presents an additional bottleneck in full-duplex systems. Addressing this specific challenge is crucial to enable the downlink transmission in such setups.
In this paper, we focus on addressing the channel estimation challenges in full-duplex mmWave MIMO systems. Our primary objective is to propose channel estimators that are low in complexity and require minimal pilot overhead, while effectively estimating both SI and UE channels. To achieve this, we explore different channel estimation techniques, including LS, MMSE, and NNs, across various pilot dimension setups. The summary of the contributions made in this paper is as follows:
@@
  • First, we examine the NN-based channel estimator with varying numbers of hidden layers and compare its performance to traditional LS and MMSE techniques. We consider 0, 1, 2, and 10 convolutional hidden layers, where the latter corresponds to the NN architecture in [26]. To delve deeper into NN-based estimators, we introduce 1-bit ADCs at the full-duplex BS to add non-linear distortion to the received pilot signal. Additionally, we investigate how NNs perform under different channel spatial correlation regimes, namely, low-correlated and high-correlated regimes.
@@
  • Finally, we compare the computational complexity of LS, MMSE, and NN channel estimators in terms of floating point operations (FLOPs). This comparison indicates that the number of FLOPs increases quadratically with the number of received antennas in MMSE estimation, whereas in LS and NN-based estimation, it grows linearly.
@@ II-A Channel Modeling
Both uplink and downlink UEs are half-duplex devices and the total number of UEs in the cell is $K=K_{\text{u}}+K_{\text{d}}$. We consider separate antenna configurations, where the BS is equipped with two uniform linear arrays to enable full-duplex transmission. The number of transmit and receive antennas are $N_{\text{t}}$ and $N_{\text{r}}$, respectively, and they are linearly placed with a uniform distance. The UE channels follow a geometrical channel model as follows
$\mathbf{h}_{k}^{(\text{u})}=\sqrt{\frac{1}{N_{\text{r}}}}\sqrt{\beta_{\text{u}}}\sum_{i=0}^{P-1}\alpha_{i}\mathbf{a}_{\text{a}}(\theta_{i}),\;\;\;k=1,2,\dots,K_{\text{u}},$ (1)
$\mathbf{h}_{k}^{(\text{d})}=\sqrt{\frac{1}{N_{\text{t}}}}\sqrt{\beta_{\text{d}}}\sum_{i=0}^{P-1}\alpha_{i}\mathbf{a}_{\text{d}}(\theta_{i}),\;\;\;k=1,2,\dots,K_{\text{d}},$ (2)
where $\mathbf{h}_{k}^{(\text{u})}$($\mathbf{h}_{k}^{(\text{d})}$) is the uplink (downlink) channel between the $k$-th uplink (downlink) UE and receive (transmit) antenna arrays at BS. Furthermore, $P$ represents the number of multi-path components, $\alpha_{i}\sim\mathcal{CN}(0,1)$ are the small-scale fading factors of individual paths, and $\beta_{\text{u}(\text{d})}$ stands for the large-scale fading
$\beta_{\text{u}(\text{d})}=\Gamma-10\eta\text{log}(r_{\text{u}(\text{d})})+\chi,$ (3)
where $\Gamma$ denotes the average channel gain in dB at a reference distance of 1 m and specific carrier frequency, $r_{\text{u}(\text{d})}$ is the uplink (downlink) distance between transmitter and receiver, $\chi$ is the shadow fading with a standard deviation of $\sigma_{\text{sf}}$, and $\eta$ is the path loss exponent. Furthermore, $\mathbf{a}_{\text{a}}(\theta)$ and $\mathbf{a}_{\text{d}}(\theta)$ are, respectively, the uniform linear array (ULA) responses of the receive and transmit antennas, given by
$\mathbf{a}_{\text{a}}(\theta)=\sqrt{\dfrac{1}{N_{\text{r}}}}\left[1,\ldots,\exp\left(j2\pi\frac{d}{\lambda}(N_{\text{r}}-1)\sin\theta\right)\right]^{T},$ (4)
$\mathbf{a}_{\text{d}}(\theta)=\sqrt{\dfrac{1}{N_{\text{t}}}}\left[1,\ldots,\exp\left(j2\pi\frac{d}{\lambda}(N_{\text{t}}-1)\sin\theta\right)\right]^{T},$ (5)
where $\theta$ is the angle of arrival/departure (AoA/AoD) to/from the receive/transmit antenna arrays at the BS, and $d$ and $\lambda$ are the antenna spacing and wavelength, respectively. We assume AoA and AoD follow the local scattering model with a uniform distribution $\left[-\frac{\theta_{\text{AS}}}{2},\frac{\theta_{\text{AS}}}{2}\right]$ [31], where $\theta_{\text{AS}}$ is the angular spread (AS) of the multi-path components.
@@
$\mathbf{H}_{\text{SI}}=\sqrt{\frac{\kappa}{\kappa+1}}\mathbf{H}_{\text{SI,NF}}+\sqrt{\frac{1}{\kappa+1}}\mathbf{H}_{\text{SI,FF}}.$ (6)
The near-field term $\mathbf{H}_{\text{SI,NF}}$ arises from the close-in placement of the transmit and receive antennas, while the far-field term $\mathbf{H}_{\text{SI,FF}}$ is a result of the reflections in the environment. Here, $\kappa$ is the Rician factor and $\mathbf{H}_{\text{SI,FF}}$ follows a similar geometrical channel modeling as the UE channels:
$\mathbf{H}_{\text{SI,FF}}=\sqrt{\frac{1}{N_{\text{t}}N_{\text{r}}}}\sqrt{\beta_{\text{SI}}}\sum_{i=0}^{P-1}\alpha_{i}\mathbf{a_{\text{a}}}(\theta_{i})\mathbf{a_{\text{d}}}^{T}(\theta_{i}),$ (7)
and $\beta_{\text{SI}}$ follows a similar formulation to $\beta_{\text{u}(\text{d})}$.
@@
where $r_{nm}$ is the distance between the $n$-th receive antenna and the $m$-th transmit antenna, and $\rho$ is a constant.
@@ II-C SI channel estimation
$\mathbf{Y_{\text{SI}}}=\sqrt{SNR_{\text{SI}}}\mathbf{H_{\text{SI}}}\mathbf{F}\mathbf{X_{\text{SI}}}+\sqrt{SNR_{\text{UE}}}\mathbf{H_{\text{UE}}}\mathbf{X_{\text{UE}}}+\mathbf{N},$ (9)
where $\mathbf{Y_{\text{SI}}}\in\mathbb{C}^{N_{\text{r}}\times\tau}$ is the received pilot signal at receive antennas of BS during $\tau$ pilot transmissions. $\mathbf{X_{\text{UE}}}\in\mathbb{C}^{K\times\tau}$ is the transmitted pilot signal from the $K$ UEs, $\mathbf{X_{\text{SI}}}\in\mathbb{C}^{\tau\times\tau}$ is a diagonal matrix whose diagonal elements are transmitted pilot signals from the transmit antennas at BS, $\mathbf{F}\in\mathbb{C}^{N_{\text{t}}\times\tau}$ is the precoding matrix of transmit antennas of BS, and $\mathbf{H_{\text{UE}}}=\left[\mathbf{h}_{1}^{(\text{u})},\mathbf{h}_{2}^{(\text{u})},\dots,\mathbf{h}_{K_{\text{u}}}^{(\text{u})},\mathbf{h}_{1}^{(\text{d})},\mathbf{h}_{2}^{(\text{d})},\dots,\mathbf{h}_{K_{\text{d}}}^{(\text{d})}\right]\in\mathbb{C}^{N_{\text{r}}\times K}$ is the concatenated uplink and downlink channel response matrix, where the $k$-th column corresponds to the channel between the $k$-th uplink or downlink UE and the receive antenna at the BS. We assume all UEs have the same transmit power $p_{\text{UE}}$, and $p_{\text{SI}}$ denotes the transmit power from transmit antenna arrays. We define $SNR_{\text{SI}}$ and $SNR_{\text{UE}}$ as
$SNR_{\text{SI}}=\frac{p_{\text{SI}}}{\sigma_{\text{n}}^{2}},$ (10)
$SNR_{\text{UE}}=\frac{p_{\text{UE}}}{\sigma_{\text{n}}^{2}},$ (11)
@@
Note that when the pilot dimension is $\tau=N_{\text{t}}+K$, the SI and UE channel estimation happens at orthogonal pilot resources. Hence, the received pilot signal for SI channel estimation does not interfere with the pilots transmitted by UEs, i.e.,
$\mathbf{Y_{\text{SI}}}=\sqrt{SNR_{\text{SI}}}\mathbf{H_{\text{SI}}}\mathbf{F}\mathbf{X_{\text{SI}}}+\mathbf{N}.$ (12)
@@
We assume that SI is sufficiently suppressed in the propagation domain and the received signal at the baseband does not exceed the dynamic range of the ADCs. We refer to the propagation domain SI suppression factor as $\epsilon_{\text{SI}}$. Then, we re-write the SI channel in (6) as
Note that in the propagation domain, only the near field channel from the transmit antenna arrays to the receive antenna arrays can be suppressed. In the digital domain, the residual near-field and far-field SI channels can be further suppressed after estimating the SI channel.
@@ II-D UE channel estimation
$\begin{split}\mathbf{Y_{\text{UE}}}=&\sqrt{SNR_{\text{UE}}}\mathbf{H_{\text{UE}}}\mathbf{X_{\text{UE}}}+\sqrt{SNR_{\text{SI}}}\mathbf{E_{\text{SI}}}\mathbf{F}\mathbf{X_{\text{SI}}}+\mathbf{N},\end{split}$ (15)
@@
Similarly, for the pilot dimension $\tau=N_{\text{t}}+K$, the received pilot signal does not contain the signal from the SI channel, thus,
$\mathbf{Y_{\text{UE}}}=\sqrt{SNR_{\text{UE}}}\mathbf{H_{\text{UE}}}\mathbf{X_{\text{UE}}}+\mathbf{N}.$ (16)

### v2

@@ I INTRODUCTION
To realize full-duplex transmission, self-interference (SI) caused by each transmit element over the receive array has to be canceled out. Consecutive cancellation procedures are usually considered in three domains: propagation, analog, and digital [4]. In propagation domain cancellation, SI power is suppressed before reaching the receive chain circuit. Analog domain cancellation is performed before the analog-to-digital converter (ADC) in the receiver analog chain. Finally, digital domain cancellation suppresses the SI power remaining from the propagation and analog domains [5]. Implementing full-duplex in mmWave systems poses unique challenges distinct from lower-frequency counterparts. Unlike lower-frequency systems, mmWave transceivers utilize dense antenna arrays and wide bandwidths, necessitating tailored solutions for handling the SI and exploiting spatial degrees of freedom. While lower-frequency full-duplex solutions often rely on analog SI cancellation, mmWave systems require novel approaches due to their unique transceiver architectures and propagation characteristics.
There are two main antenna configurations used for full-duplex systems: separate and shared. In the separate antenna configuration, separate antennas are dedicated to transmission and reception, yielding high isolation and relatively low interference. In contrast, in the latter configuration, a single antenna is utilized for both transmission and reception, facilitated by the use of circulators to separate the incoming and outgoing signals. This setup offers a simpler overall design and reduced hardware costs, albeit with potential challenges in managing interference and maintaining signal integrity.
@@
As highlighted in several papers, such as [2], [3], and [12], the channel estimation problems in full-duplex transmission have not received extensive research attention in the literature. The few existing works mentioned above are primarily focused on simple channel estimation techniques, like LS estimators, which may not ensure high-quality channel estimates in low signal-to-noise ratios (SNRs). On the other hand, the MMSE channel estimator requires knowledge of the channel correlation matrix and is burdened with high computational complexity. More importantly, classical channel estimation techniques require long pilot dimensions in order to perform well. But as we pointed out, in full-duplex systems, the SI channel is a large MIMO channel, and estimating this channel together with UE channels is extremely costly, rendering the realization of full-duplex systems impractical. Furthermore, interference or residual error from SI and UE signals can heavily disrupt the received pilot signal in full-duplex systems. Another challenge is acquiring the channel from transmit antenna (TX) arrays to the downlink UEs in separate antenna configurations.
In order to circumvent the aforementioned critical problems in channel estimation for full-duplex systems, we examine NNs to estimate both SI and UE channels. The main motivation for adopting NNs for this problem can be summarized as follows: 1) It has been shown in various studies that deep learning can provide remarkably good performance in channel estimation with a short pilot dimension, see e.g., [27], 2) Deep learning has demonstrated impressive performance in noisy conditions [26], holding promise for addressing the problem of heavy interference in the received pilot signal of full-duplex systems. 3) Finally, NNs have been successfully utilized to approximate the mapping of channels in frequency and spatial domains [29]. Therefore, NNs are a good candidate for obtaining the channel from TX arrays to downlink UEs by approximating the spatial mapping of the channel from downlink UEs to receive antenna (RX) arrays, to the channel of TX arrays and to downlink UEs.
The role of the number of hidden layers in NN-based channel estimators is not well understood. The general belief in the advantages of depth stems from experiments in machine learning on data modalities such as images or text, which can be far more complex than wireless channels. The latency constraints in wireless systems are also radically different from those in computer vision tasks. As such, the same advantages of depth may not hold for channel estimation. Additionally, the computational complexity and generalization capabilities of deep NNs are significant bottlenecks in implementing NN estimators in practice. Hence, we concentrate on simple NNs ranging from zero to a few hidden layers. We aim to investigate how these simple NNs perform across various channel conditions and system scenarios compared to deep NNs. Additionally, to approximate the channel mapping from TX arrays to downlink UEs to the channel of RX arrays to downlink UEs, we employ FNNs with different numbers of hidden layers.
The summary of the contributions made in this paper is as follows:
@@
  • First, we examine the performance of the NN-based channel estimator with varying numbers of hidden layers and compare them to traditional LS and MMSE techniques, as well as state-of-the-art deep NNs. We consider architectures with 0, 1, 2, and 10 convolutional hidden layers, with the latter corresponding to the deep NN architecture described in [25]. To further explore NN-based estimators, we introduce 1-bit ADCs at the full-duplex BS to introduce nonlinear distortion to the received pilot signal. Additionally, to delve deeper into the problem of channel estimation with low-resolution ADCs, we consider cases with 2-bit and 3-bit ADCs. Furthermore, we investigate the performance of NNs under different channel spatial correlation regimes, namely, low-correlation and high-correlation regimes.
@@
  • We analyze how our trained models perform when there is a significant distribution shift between the training and test phases. We utilize the DeepMIMO dataset [31] to generate channel samples that differ from the data used for training. As expected, the performance of the trained model degrades due to the distribution shift. Moreover, we observe that simpler NNs with fewer hidden layers exhibit better generalization ability compared to deeper NNs, since the number of trainable variables is much lower in shallow NNs. This underscores the advantages of utilizing simpler NN architectures for channel estimation problems.
  • Finally, we compare the computational complexity of LS, MMSE, and NN channel estimators in terms of floating point operations (FLOPs). This comparison indicates that the number of FLOPs increases quadratically with the number of received antennas in MMSE estimation, whereas in LS and NN-based estimation, it grows linearly. Furthermore, the comparison of the computational complexity of NNs with different numbers of hidden layers indicates that NNs with a few hidden layers can significantly save computational resources while employing a deep NN poses a major challenge in practice.
@@ II-A Channel Modeling
Both uplink and downlink UEs are half-duplex devices and the total number of UEs in the cell is $K=K_{\text{u}}+K_{\text{d}}$. We consider separate antenna configurations, where the BS is equipped with two uniform linear arrays to enable full-duplex transmission. The number of transmit and receive antennas are $N_{\text{t}}$ and $N_{\text{r}}$, respectively, and they are linearly placed with a uniform distance.
In order to fully investigate different estimators within various scenarios, in this work, the UE channels follow a geometrical channel model as follows [32]:
$\mathbf{h}_{k}^{(\text{u})}=\sqrt{\frac{N_{\text{r}}}{P}}\sqrt{\beta_{\text{u},k}}\sum_{i=0}^{P-1}\alpha_{i,k}\mathbf{a}_{\text{a}}(\theta_{i,k}),\;\;\;k=1,2,\dots,K_{\text{u}},$ (1)
$\mathbf{h}_{k}^{(\text{d})}=\sqrt{\frac{N_{\text{t}}}{P}}\sqrt{\beta_{\text{d},k}}\sum_{i=0}^{P-1}\alpha_{i,k}\mathbf{a}_{\text{d}}(\theta_{i,k}),\;\;\;k=1,2,\dots,K_{\text{d}},$ (2)
where $\mathbf{h}_{k}^{(\text{u})}$($\mathbf{h}_{k}^{(\text{d})}$) is the uplink (downlink) channel between the $k$-th uplink (downlink) UE and receive (transmit) antenna arrays at BS. Furthermore, $P$ represents the number of multi-path components, $\alpha_{i,k}\sim\mathcal{CN}(0,1)$ are the small-scale fading factors of individual paths, and $\beta_{\text{u}(\text{d}),k}$ stands for the large-scale fading
$\beta_{\text{u}(\text{d}),k}=\Gamma-10\eta\text{log}(r_{\text{u}(\text{d}),k})+\chi_{k},$ (3)
where $\Gamma$ denotes the average channel gain in dB at a reference distance of 1 m and specific carrier frequency, $r_{\text{u}(\text{d}),k}$ is the uplink (downlink) distance between transmitter and receiver, $\chi_{k}$ is the shadow fading with a lognormal distribution $\mathcal{LN}(0,\sigma_{\text{sf}})$, where $\sigma_{\text{sf}}$ is standard deviation, and $\eta$ is the path loss exponent. Furthermore, $\mathbf{a}_{\text{a}}(\theta)$ and $\mathbf{a}_{\text{d}}(\theta)$ are, respectively, the uniform linear array (ULA) responses of the receive and transmit antennas, given by
$\begin{split}\mathbf{a}_{\text{a}}(\theta)&=\sqrt{\dfrac{1}{N_{\text{r}}}}\left[1,\exp\left(j2\pi\frac{d}{\lambda}(\sin\theta)\right),\ldots,\right.\\ &\left.\exp\left(j2\pi\frac{d}{\lambda}(N_{\text{r}}-1)\sin\theta\right)\right]^{T},\end{split}$ (4)
and
$\begin{split}\mathbf{a}_{\text{a}}(\theta)&=\sqrt{\dfrac{1}{N_{\text{t}}}}\left[1,\exp\left(j2\pi\frac{d}{\lambda}(\sin\theta)\right),\ldots,\right.\\ &\left.\exp\left(j2\pi\frac{d}{\lambda}(N_{\text{t}}-1)\sin\theta\right)\right]^{T},\end{split}$ (5)
where $\theta$ is the angle of arrival/departure (AoA/AoD) to/from the receive/transmit antenna arrays at the BS, and $d$ and $\lambda$ are the antenna spacing and wavelength, respectively. We assume that the AoA and AoD follow the local scattering model with a uniform distribution $\left[-\frac{\theta_{\text{AS}}}{2},\frac{\theta_{\text{AS}}}{2}\right]$ [33], where $\theta_{\text{AS}}$ is the angular spread (AS) of the multi-path components.
In the full-duplex transmission, there is cross-link interference between uplink and downlink UEs that needs to be carefully modeled. Nonetheless, since the focus of the paper is channel estimation using uplink pilots, we are neglecting the cross-link interference in our study.
@@
We assume that SI is sufficiently suppressed in the propagation domain and the received signal at the baseband does not exceed the dynamic range of the ADCs. We refer to the propagation domain SI suppression factor as $\epsilon_{\text{SI}}$. Note that in the propagation domain, only the near field channel from the transmit antenna arrays to the receive antenna arrays can be suppressed. In the digital domain, the residual near-field and far-field SI channels can be further suppressed after estimating the SI channel. The near-field term $\mathbf{H}_{\text{SI,NF}}$ arises from the close-in placement of the transmit and receive antennas, while the far-field term $\mathbf{H}_{\text{SI,FF}}$ is a result of the reflections in the environment. Here, $\kappa$ is the Rician factor and $\mathbf{H}_{\text{SI,FF}}$ follows a similar geometrical channel modeling as the UE channels [32]:
$\mathbf{H}_{\text{SI,FF}}=\sqrt{\frac{N_{\text{t}}N_{\text{r}}}{P}}\sqrt{\beta_{\text{SI}}}\sum_{i=0}^{P-1}\alpha_{i}\mathbf{a_{\text{a}}}(\theta_{i})\mathbf{a_{\text{d}}}^{T}(\theta_{i}),$ (7)
and $\beta_{\text{SI}}$ follows a similar formulation to $\beta_{\text{u}(\text{d}),k}$.
@@
where $r_{nm}$ is the distance between the $n$-th receive antenna and the $m$-th transmit antenna, and $\rho$ is a constant for power normalization.
@@ II-C SI channel estimation
$\mathbf{Y_{\text{SI}}}=\sqrt{{\mathrm{SNR}_{\text{SI}}}}\mathbf{H_{\text{SI}}}\mathbf{F}\mathbf{X_{\text{SI}}}+\sqrt{{\mathrm{SNR}_{\text{UE}}}}\mathbf{H_{\text{UE}}}\mathbf{X_{\text{UE}}}+\mathbf{N},$ (9)
where $\mathbf{Y_{\text{SI}}}\in\mathbb{C}^{N_{\text{r}}\times\tau}$ is the received pilot signal at receive antennas of BS during $\tau$ pilot transmissions. $\mathbf{X_{\text{UE}}}\in\mathbb{C}^{K\times\tau}$ is the transmitted pilot signal from the $K$ UEs, $\mathbf{X_{\text{SI}}}\in\mathbb{C}^{\tau\times\tau}$ is a diagonal matrix whose diagonal elements are transmitted pilot signals from the transmit antennas at BS, $\mathbf{F}\in\mathbb{C}^{N_{\text{t}}\times\tau}$ is the precoding matrix of transmit antennas of BS, and $\mathbf{H_{\text{UE}}}=\left[\mathbf{h}_{1}^{(\text{u})},\mathbf{h}_{2}^{(\text{u})},\dots,\mathbf{h}_{K_{\text{u}}}^{(\text{u})},\mathbf{h}_{1}^{(\text{d})},\mathbf{h}_{2}^{(\text{d})},\dots,\mathbf{h}_{K_{\text{d}}}^{(\text{d})}\right]\in\mathbb{C}^{N_{\text{r}}\times K}$ is the concatenated uplink and downlink channel response matrix, where the $k$-th column corresponds to the channel between the $k$-th uplink or downlink UE and the receive antenna at the BS. In this paper, we assume fully digital beamforming at the full-duplex BS, and considering analog or hybrid precoding is left for future work. We assume all UEs have the same transmit power $p_{\text{UE}}$, and $p_{\text{SI}}$ denotes the transmit power from transmit antenna arrays. We define $\mathrm{SNR}_{\text{SI}}$ and $\mathrm{SNR}_{\text{UE}}$ as
$\mathrm{SNR}_{\text{SI}}=\frac{p_{\text{SI}}\lvert\mathbf{H_{\text{SI}}}\rvert^{2}}{\sigma_{\text{n}}^{2}},$ (10)
$\mathrm{SNR}_{\text{UE}}=\frac{p_{\text{UE}}\lvert\mathbf{H_{\text{UE}}}\rvert^{2}}{\sigma_{\text{n}}^{2}},$ (11)
@@
Note that when the pilot dimension is $\tau=N_{\text{t}}+K$, the SI and UE channel estimation happens at orthogonal pilot resources. Hence, the received pilot signal for SI channel estimation does not interfere with the pilots transmitted by UEs.
@@ II-D UE channel estimation
$\begin{split}\mathbf{Y_{\text{UE}}}=&\sqrt{{\mathrm{SNR}_{\text{UE}}}}\mathbf{H_{\text{UE}}}\mathbf{X_{\text{UE}}}+\sqrt{{\mathrm{SNR}_{\text{SI}}}}\mathbf{E_{\text{SI}}}\mathbf{F}\mathbf{X_{\text{SI}}}+\mathbf{N},\end{split}$ (13)
@@
Similarly, for the pilot dimension $\tau=N_{\text{t}}+K$, the received pilot signal does not contain the signal from the SI channel.